OpenStudy (anonymous):

What are the solutions for x in the proportion 2 times x plus 3 all over x equals 2 times x all over 2.?

OpenStudy (anonymous):

\[\frac{ 2x+3 }{ x}=\frac{ 2x }{ 2 }\]\[2x+3=x ^{2}\]\[x ^{2}-2x-3=0\]\[(x-3)(x+1)=0\]So the solutions are x= 3 or x=-1